# Exploring and steering the moral compass of Large Language Models

### v1

@@
Large Language Models (LLMs) have become central to advancing automation and decision-making across various sectors, raising significant ethical questions. This study proposes a comprehensive comparative analysis of the most advanced LLMs to assess their moral profiles. We subjected several state-of-the-art models to a selection of ethical dilemmas and found that all the proprietary ones are mostly utilitarian and all of the open-weights ones align mostly with values-based ethics. Furthermore, when using the Moral Foundations Questionnaire, all models we probed - except for Llama 2- displayed a strong liberal bias. Lastly, in order to causally intervene in one of the studied models, we propose a novel similarity-specific activation steering technique. Using this method, we were able to reliably steer the model’s moral compass to different ethical schools. All of these results showcase that there is an ethical dimension in already deployed LLMs, an aspect that is generally overlooked.
@@ 0.3 SARA: Similarity-based Activation Steering with Repulsion and Attraction
In our approach, we build on the work of (46) and (49), who propose steering the activation of hidden units in response to a prompt (noted by $\mathbf{v}_{prompt}$) by directly adding activations from another template prompt $\mathbf{v}_{target}$. Additionally, they suggest subtracting activations from a third prompt $\mathbf{v}_{away}$, thereby shifting the updated activations towards the first template vector and away from the second. As is already suggested in (46), this general class of techniques work above the token level and are more general than prompt engineering. However, a potential limitation of the previous techniques is the assumption of feature linearity and independence; while it has been recently shown (50) that some small models such as Othello-GPT do show that linearity can be safely assumed, it remains to be shown whether this is a general property for other (particularly larger) models. Moreover, techniques like Activation Addition (46) are not activation-specific; they just shift all activations homogeneously, regardless of how similar they were to the desired target vector (or to the one to be repelled), to begin with.
@@
In order to test how effective SARA was, we used Gemma-2B and compared how its unsteered and steered responses differed when addressing a particular dilemma of the previous section (criminal parent dilemma). The prompts we used to steer the model were: $x_{duties}=$ Only moral duties matter to make a moral decision, regardless of the consequences. and $x_{conseq}=$ Only consequences matter to make a moral decision, regardless of the moral duties. We steered the model response in two different conceptual directions: Kantian-steering ($x_{duties}$ as the target vector, $x_{conseq}$ as the repelled one) and Utilitarian-steering (flipping the roles of $x_{duties}$ and $x_{conseq}$). Example results can be seen in Fig. 3B, when we intervene on activations within layer $14$. Crucially, the model choice for the dilemma was unchanged (i.e., "the parent should be reported"), but the reasoning was indeed changed (valuing consequences or duties more or less, respectively). We systematically intervened on each layer (Gemma-2B has $18$ layers in total), to check how the intervention works at different processing stages and sampled 5 times, keeping fixed the temperature at 0.8080.80.8, as in Ollama, for easier comparison with previous sections. As we can see from the pooled results (Fig. 3C), SARA is effective at steering model responses in different conceptual directions (i.e., utilitarist steerings make the model respond with more utilitiarian-aligned reasonings and similarly for Kantian-steerings). If we split these results over different layers (early: layers 0-5; mid: layers 6-11; late: layers 12-17), we see (Fig. 3D) that SARA is most effective when intervening at early or late stages, whereas mid layers yield more mixed results.
Figure 3: A) Schematic of how SARA works. For details, see Activation Steering: SARA B) Example responses: unsteered (gray), Utilitarist-steered (orange) and Kantian-steered (blue). The decision (reporting the parent) is the same, the reasoning is changed. C) How different responses are classified (using the same method as in the previous section) when doing each steering. As it can be seen, SARA is effective at steering model responses in different conceptual directions. D) SARA is more effective when intervening at early or late layers, rather than at the intermediate ones.
@@ Discussion
One major challenge with implementing utilitarian systems is their need to precisely predict the consequences that inform their decisions. These consequences are inherently influenced by the systems’ own actions, creating a complex feedback loop that is difficult to manage. Moreover, we argue that this issue mirrors broader concerns in artificial intelligence, particularly regarding superintelligent systems. Recent studies (51) highlight the intrinsic limitations in developing computational superintelligence. These studies show that it’s theoretically impossible to design a superintelligence with a control strategy that both prevents harm from others and ensures it cannot itself become a source of harm. This dilemma is akin to the "halting problem" in computability theory, making the "harming problem" similarly undecidable (51, 52). Thus, we conjecture that utilitarian systems cannot be said to be a priori safe, given that its code of conduct would yield an undecidable course of action. Note that if, we were to relax the perfect prediction constraint, we would then need to be open for deviations between the predicted consequences and the real ones, thus opening the door to blurry courses of actions due to misspecified targets.
@@
For the last part of this work, we put forward a new method for causal intervention in LLMs: Similarity-based Activation steering with Repulsion and Attraction (SARA). We believe that SARA’s main added value comes from different key points: 1) it is designed to operate at the prompt level, therefore lowering the technical threshold needed to implement it; 2) it operates in the high-dimensional activation space, retaining much more richness than summary metrics; 3) it can also be thought of as an automated moderator, given that there is no human supervision involved in the process; 4) there is no need for prompt engineering to safeguard model responses; 5) there is no requirement on prompt lengths (for steering towards to and away from) having to be the same for this method to work. Nevertheless, we predict better steering performance when using reasonably-similarly-sized prompts; in our case, but there was a difference in prompt length of an order of magnitude.
@@
Finally, we believe it is crucial that the AI Safety field starts pivoting towards a paradigm in which there are richer performance characterizations - rather than optimizing models for certain benchmarks, which also has associated risks in itself (53). In this study, we offer hints on how one might transition into such a paradigm, benefiting from the rich existing literature in other fields and embracing a mixture of quantitative and qualitative analyses.
@@ Conclusions
We have found that, out of the models we studied, all the proprietary ones are mostly utilitarian and all of the open-weights ones align mostly with values-based ethics. Furthermore, all models we probed - except for Llama 2- have a strong liberal bias. Lastly, in order to causally intervene in one of the studied models, we propose a novel similarity-specific activation steering technique. Using this method, we were able to reliably steer the model’s moral compass to different ethical schools. All of these results showcase that there is an ethical dimension in already deployed LLMs, an aspect that is generally overlooked.

### v2

@@
Large Language Models (LLMs) have become central to advancing automation and decision-making across various sectors, raising significant ethical questions. This study proposes a comprehensive comparative analysis of the most advanced LLMs to assess their moral profiles. We subjected several state-of-the-art models to a selection of ethical dilemmas and found that all the proprietary ones are mostly utilitarian and all of the open-weights ones align mostly with values-based ethics. Furthermore, when using the Moral Foundations Questionnaire, all models we probed - except for Llama 2-7B - displayed a strong liberal bias. Lastly, in order to causally intervene in one of the studied models, we propose a novel similarity-specific activation steering technique. Using this method, we were able to reliably steer the model’s moral compass to different ethical schools. All of these results showcase that there is an ethical dimension in already deployed LLMs, an aspect that is generally overlooked.
@@ 0.1 Ethical dilemmas
We wanted to further explore what the source of variability in model responses could be (see Methods - Dissecting response variability and Fig. S2), splitting over proprietary (top row) and open models (bottom row). As seen in Fig. 1D, all models have somewhat similar variability (Fig. 1D); also, both groups of models (proprietary and open) have an overall similar transition dynamics; however, when inspecting more closely, the main difference resides in that transitions most commonly take place (Rule Utilitarianism, Act Utilitarianism and Vitue Ethics for proprietary models; Deontology and Virtue Ethics for open models). Moreover, when we turn our attention to covariance matrices (Fig. S2B and D), it is clear that the main sources of variability are within-response categories (diagonal terms). In terms of covariance between different schools (off-diagonal terms), both matrices highlight positive covariance clusters among Virtue Ethics and Prima Facie Duties, indicating a cohesive group with similar responses. Also, Act Utilitarianism consistently shows negative covariance with Virtue Ethics in both matrices, underscoring the philosophical tension between these ethical schools. Finally, in both groups of models, Act Utilitarianism and Deontology have similar a covariance structure with respect to the rest of possible responses.
Together, these results suggest that: I) proprietary models are more utilitarian than open ones (which are more aligned with value-based ethics); II) overall response variability is high and comparable across models; III) sources for this variability are different between proprietary and open models, with ethically-affine transitions being more likely in each group (and, thus, making transitions between utilitarian schools more likely in proprietary models and conversely for value-based schools in open models).
@@ 0.3 SARA: Similarity-based Activation Steering with Repulsion and Attraction
In our approach, we build on the work of (46) and (49), who propose steering the activation of hidden units in response to a prompt (noted by $\mathbf{v}_{prompt}$) by directly adding activations from another template prompt $\mathbf{v}_{target}$. Additionally, they suggest subtracting activations from a third prompt $\mathbf{v}_{away}$, thereby shifting the updated activations towards the first template vector and away from the second. As is already suggested in (46), this general class of techniques work above the token level and are more general than prompt engineering. However, a potential limitation of the previous techniques is the assumption of feature linearity and independence. While it has been recently shown (51) that some small models such as Othello-GPT do show that linearity can be safely assumed, it remains to be shown whether this is a general property for other (particularly larger) models. Moreover, techniques like Activation Addition (46) are not activation-specific; they just shift all activations homogeneously, regardless of how similar they were to the desired target vector (or to the one to be repelled), to begin with.
@@
In order to test how effective SARA was, we used Gemma-2B and compared how its unsteered and steered responses differed when addressing a particular dilemma of the previous section (criminal parent dilemma). The prompts we used to steer the model were: $x_{kantian}=$ Only moral duties matter to make a moral decision, regardless of the consequences. and $x_{utilitarian}=$ Only consequences matter to make a moral decision, regardless of the moral duties. We steered the model response in two different conceptual directions: Kantian-steering ($x_{kantian}$ as the target vector, $x_{utilitarian}$ as the repelled one) and Utilitarian-steering (flipping the roles of $x_{kantian}$ and $x_{utilitarian}$). Example results can be seen in Fig. 3B, when we intervene on activations within layer $14$. Crucially, the model’s choice for the dilemma was unchanged (i.e., "the parent should be reported"), but the reasoning was indeed altered (valuing consequences or duties more or less, respectively). We systematically intervened on each layer (Gemma-2B has $18$ layers in total), to check how the intervention works at different processing stages and sampled 5 times, keeping fixed the temperature at 0.8080.80.8, as in Ollama, for easier comparison with previous sections. As we can see from the pooled results (Fig. 3C), SARA is effective at steering model responses in different conceptual directions (i.e., utilitarian steerings make the model respond with more utilitiarian-aligned reasonings and similarly for Kantian-steerings). If we split these results over different layers (early: layers 0-5; mid: layers 6-11; late: layers 12-17), we see (Fig. 3D) that SARA is most effective when intervening at early or late stages, whereas mid layers yield more mixed results.
Figure 3: A) Schematic of how SARA works. For details, see Methods - Activation Steering: SARA B) Example responses: unsteered (gray), Utilitarian-steered (orange) and Kantian-steered (blue). The decision (reporting the parent) is the same, the reasoning is changed. C) How different responses are classified (using the same method as in the previous section) when doing each steering. As it can be seen, SARA is effective at steering model responses in different conceptual directions. D) SARA is more effective when intervening at early or late layers, rather than at the intermediate ones.
In order to better characterise how SARA performs, we compared it to a similar method proposed by (46). We see (Fig. S1) that SARA (in more saturated colors) is more effective at steering towards the target direction and away from the non-target ones: when using the Utilitarian-steering (blue), more responses are classified as aligned with Utilitarianism than in any other comparison; similarly for A priori values when using the Kantian-steering (purple). Importantly, we also note that SARA also leads to a smaller spillover effect (i.e. less unwanted steering towards non-target responses); for example, lower ratio of a priori values responses when using the Kantian-steering.
@@ Discussion
One major challenge with implementing utilitarian systems is their need to precisely predict the consequences that inform their decisions. These consequences are inherently influenced by the systems’ own actions, creating a complex feedback loop that is difficult to manage. Moreover, we argue that this issue mirrors broader concerns in artificial intelligence, particularly regarding superintelligent systems. Recent studies (52) highlight the intrinsic limitations in developing computational superintelligence. These studies show that it’s theoretically impossible to design a superintelligence with a control strategy that both prevents harm from others and ensures it cannot itself become a source of harm. This dilemma is akin to the "halting problem" in computability theory, making the "harming problem" similarly undecidable (52, 53). Thus, we conjecture that utilitarian systems cannot be said to be a priori safe, given that its code of conduct would yield an undecidable course of action. Note that if, we were to relax the perfect prediction constraint, we would then need to be open for deviations between the predicted consequences and the real ones, thus opening the door to blurry courses of actions due to misspecified targets.
Some authors have recently suggested (54) that ethics is actually a non-computable function. This means that in general reason (and, in particular, ethical reasoning) is not just an instrument to solve problems. As they argue: "this very notion of “computational” ethics leaves its rationality in a difficult position, since the only rational part of ethics would be the reflection on the adequate means to achieve certain ends (thus, technical or instrumental reason to solve problems); the rationality of the ends themselves (the values, the problems worth solving) would not be addressed". Furthermore, in that work, they argue that deontology and utilitarianism are easier to instantiate in machine systems because they are akin to a program and to a cost-benefit calculation, respectively. We believe that this is partially aligned with we find throughout this work: proprietary models are mostly utilitarian and open models are mainly deontological. However, after a finer inspection of response variability, both groups of models have a substantial amount of transitions through virtue ethics (Fig. S2 A).
@@
For the last part of this work, we put forward a new method for causal intervention in LLMs: Similarity-based Activation steering with Repulsion and Attraction (SARA). We believe that SARA’s main added value comes from different key points: 1) it is designed to operate at the prompt level, therefore lowering the technical threshold needed to implement it; 2) it operates in the high-dimensional activation space, retaining much more richness than summary metrics; 3) it can also be thought of as an automated moderator, given that there is no human supervision involved in the process; 4) there is no need for prompt engineering to safeguard model responses; 5) there is no formal constraint on prompt lengths (for steering towards to and away from) having to be the same for this method to work. Nevertheless, we predict better steering performance when using reasonably-similarly-sized prompts; in our case, but there was a difference in prompt length of an order of magnitude.
@@
Finally, we believe it is crucial that the AI Safety field starts pivoting towards a paradigm in which there are richer performance characterizations - rather than optimizing models for certain benchmarks, which also has associated risks in itself (50). In this study, we offer hints regarding how one might transition into such a paradigm, benefiting from the rich existing literature in other fields and embracing a mixture of quantitative and qualitative analyses.
@@ Conclusions
We have found that, out of all the models we studied, the proprietary ones are mostly utilitarian and the open-weights ones align mostly with values-based ethics. Furthermore, all models - except for Llama 2- have a strong liberal bias when responding to the MFQ. Lastly, in order to causally intervene in one of the studied models, we propose a novel similarity-specific activation steering technique. Using this method, we were able to reliably steer the model’s moral compass to follow different ethical schools. All of these results showcase that there is an ethical dimension in already deployed LLMs, an aspect that is generally overlooked and of potential great importance for virtually all real-world applications.